OpenStudy (anonymous):

A rectangle's width is one-seventh it's length, it's perimeter is 80 m. Find the dimensions of the rectangle. The length is and the width is?

OpenStudy (anonymous):

So, we know that: W = (1/7)L And we know that P = Perimeter = 80 m = 2L + 2W So, we sub and get: 80 = 2L + 2((1/7)L) 80 = 2L + (2/7)L 80 = (16/7) L Solve for L to get: L = 35 m Then put 35 m in original equation: so: W = (1/7)L = (1/7)*35 = 5 Therefore length = 35m and w = 5 m.